# Empirical evaluation of normalizing flows in Markov Chain Monte Carlo

### v1

@@ 4.1.2 Best-case analysis for specific target families
Jump HMC ranks best on each family and is tied with Jump MH on non-Gaussian unimodal targets. In gradient-free sampling, Jump MH performs best on synthetic targets, while IMH is better on real-world targets. Hoffman et al. (2019) show that NeuTra HMC explores geometrically complex targets more efficiently than HMC. However, we find their ranks to be the same on non-Gaussian targets. Due to the high uncertainty in HMC, it is plausible that some NFs allow NeuTra HMC to rank better, though this is not the case in general. Moreover, NeuTra HMC is actually worse for Gaussian targets and does not improve results in general (see Figures 1(a) and 1(b)).
@@ 4.2.1 Jump performance on different target families
NICE $-0.43\pm 0.00$.31) $1.45\pm 0.00$.14) $-0.80\pm 0.00$.25) $0.00\pm 0.00$.32) $-0.11\pm 0.00$.22)
Real NVP $-0.80\pm 0.00$.25) $-1.16\pm 0.00$.43) $0.29\pm 0.00$.36) $0.29\pm 0.00$.32) $-0.11\pm 0.00$.22)
C-LR-NSF $-0.14\pm 0.00$.72) $0.72\pm 0.00$.58) $-0.51\pm 0.00$.30) $-0.40\pm 0.00$.19) $-0.24\pm 0.00$.20)
C-RQ-NSF $-0.87\pm 0.00$.19) $1.01$ $-0.51\pm 0.00$.38) $0.43\pm 0.00$.39) $0.00\pm 0.00$.24)
C-NAFdeep $0.58\pm 0.00$.19) $-0.14\pm 0.00$.87) $0.72\pm 0.00$.35) $0.80\pm 0.00$.25) $0.63\pm 0.00$.17)
@@
i-ResNet $-0.80\pm 0.00$.36) $-0.14\pm 0.00$.58) $-0.07\pm 0.00$.60) $-0.98\pm 0.00$.25) $-0.64\pm 0.00$.20)
ResFlow $0.22\pm 0.00$.62) $-0.29\pm 0.00$.14) $0.80\pm 0.00$.18) $-0.18\pm 0.00$.33) $0.11\pm 0.00$.21)
CNFEuler $0.14\pm 0.00$.35) $-0.87\pm 0.00$.43) $-0.36\pm 0.00$.62) $-0.54\pm 0.00$.31) $-0.39\pm 0.00$.21)
CNFRK $0.29\pm 0.00$.65) $0.00\pm 0.01$.59) $-0.07\pm 0.00$.48) $-0.72\pm 0.00$.27) $-0.27\pm 0.00$.26)
CNFRK(R) $-0.43\pm 0.00$.31) $-1.16\pm 0.00$.14) $-0.80\pm 0.00$.27) $-0.54\pm 0.00$.32) $-0.64\pm 0.00$.17)
@@
The best-performing architectures across all targets are CNFRK(R), i-ResNet, and CNFEuler. All obtain $\overline{r}$ on each family except CNFEuler on Gaussians. CNFRK(R) achieves better overall $\overline{r}$ than CNFRK, suggesting that regularization is beneficial for continuous NF models. NAF models rank the worst, which could be due to having a very high parameter count. A high parameter count increases the space of possible solutions, which can result in slower training. Additionally, NAF models use neural network transformers, which are more difficult to optimize than affine or spline maps. Residual NFs and continuous NFs have fewer restrictions on the form of their bijections than autoregressive NFs. This could explain their better performance on real-world targets, whose complex geometry requires expressive bijections to be modeled. Having observed how the bijection form can contribute to sampling quality, we also investigate the effect of the NF parameter count. In Figure 2, we plot the trainable parameter count against the SR of the NF.
@@
We find a noticeable improvement in performance with $10^{3}$ to $10^{4}$ parameters where we observe a dip in SR. The results verify our claim that one must choose hyperparameters that create a suitably expressive architecture. Furthermore, they suggest that NF parameter counts in this range are a suitable default choice. For SVI with Real NVP, Agrawal and Domke (2024) generally recommend using 10 or more for targets with complex geometry, as well as many hidden units for high-dimensional targets. Our analysis shows that these recommendations do not directly translate into NFMC, as the best results are generally achieved by NFs with few parameters.
@@ 4.2.2 Preconditioning quality on different target families
C-LR-NSF $-0.62\pm 0.00$.42) $0.33\pm 0.00$.33) $-0.37\pm 0.00$.49) $0.20\pm 0.00$.32) $-0.10\pm 0.00$.21)
C-RQ-NSF $-0.87\pm 0.00$.33) $0.33\pm 0.00$.17) $-0.04\pm 0.00$.25) $-0.45\pm 0.00$.33) $-0.36\pm 0.00$.18)
C-NAFdeep $1.07\pm 0.00$.08) $0.50\pm 0.00$.17) $1.40\pm 0.00$.11) $0.62\pm 0.00$.41) $0.88\pm 0.00$.20)
C-NAFdense $0.33\pm 0.00$.45) $-0.91\pm 0.00$.58) $1.11\pm 0.00$.23) $0.56\pm 0.00$.30) $0.47\pm 0.00$.22)
C-NAFboth $0.04\pm 0.00$.32) $-0.99\pm 0.00$.33) $0.25\pm 0.00$.29) $0.08\pm 0.00$.37) $-0.02\pm 0.00$.20)
IAF $0.54\pm 0.00$.44) $1.16\pm 0.00$.33) $-0.04\pm 0.00$.60) $0.39\pm 0.00$.27) $0.41\pm 0.00$.21)
IA-LR-NSF $-0.25\pm 0.00$.55) $-0.25\pm 0.00$.08) $0.37\pm 0.00$.62) $-0.07\pm 0.00$.41) $-0.03\pm 0.00$.25)
IA-RQ-NSF $-0.62\pm 0.00$.44) $-0.50\pm 0.00$.33) $-0.87\pm 0.00$.31) $-0.30\pm 0.00$.28) $-0.52\pm 0.00$.17)
IA-NAFdeep $-0.83\pm 0.00$.24) $-1.16\pm 0.00$.33) $-0.37\pm 0.00$.40) $-0.45\pm 0.00$.38) $-0.59\pm 0.00$.20)
IA-NAFdense $-0.54\pm 0.00$.27) $-0.91\pm 0.00$.41) $-0.87\pm 0.00$.37) $-0.26\pm 0.00$.34) $-0.53\pm 0.00$.19)
IA-NAFboth $0.29\pm 0.00$.25) $-0.58\pm 0.00$.08) $0.12\pm 0.00$.46) $-0.25\pm 0.00$.38) $-0.07\pm 0.00$.20)
i-ResNet $-0.58\pm 0.00$.27) $-0.58\pm 0.00$.58) $-0.70\pm 0.00$.18) $-0.39\pm 0.00$.21) $-0.52\pm 0.00$.13)
ResFlow $-0.95\pm 0.00$.39) $-1.07\pm 0.00$.08) $-0.62\pm 0.00$.24) $-0.43\pm 0.00$.22) $-0.66\pm 0.00$.14)
Planar $0.29\pm 0.00$.55) $0.25\pm 0.00$.08) $-0.04\pm 0.00$.21) $-0.11\pm 0.00$.31) $0.04\pm 0.00$.18)
Radial $-1.16\pm 0.00$.39) $-1.32\pm 0.00$.33) $-1.57\pm 0.00$.08) $-1.07\pm 0.00$.21) $-1.23\pm 0.00$.13)
Sylvester $-0.74\pm 0.00$.20) $-0.58\pm 0.01$.07) $0.00\pm 0.00$.24) $-0.67\pm 0.00$.34) $-0.47\pm 0.00$.19)
CNFEuler $1.49$ $0.99$ $-0.00\pm 0.00$.58) $0.58\pm 0.00$.39) $0.70\pm 0.00$.24)
@@
As groups, both IA and residual NFs perform the best across all targets (last column). This is unsurprising because IA and radial flows were designed to improve variational inference (Tabak and Turner, 2013; Rezende and Mohamed, 2015; Kingma et al., 2016), which forms the crucial NF warm-up phase in NeuTra MCMC. Interestingly, the radial flow performs better than IA within NeuTra MCMC despite the fact that IAF has been proposed to address the limitations of the former. NeuTra MCMC can benefit from simpler preconditioning of the radial flow, which does not use neural networks, compared to more expressive bijections. However, the form of the bijection could play a role, as the simple planar NF ranks worse than radial and Sylvester flows.
@@
Our results also shed light on Hoffman et al. (2019), who precondition HMC with the IAF bijection, which improves moment estimates over HMC on the funnel target. However, the second column of Table 3 shows that IAF is one of the worst-performing NFs for the funnel and Rosenbrock distributions. The discrepancy could be due to differences in implementation, namely the number of chains: they use 16384 parallel chains on the GPU, whereas we use 100 on the CPU. We cannot afford such a big number of GPU chains, as it would require a prohibitive amount of computational resources for a fair comparison with other, usually slower NFs. Hoffman et al. (2019) also use a bigger SVI batch size compared to our single-sample unbiased loss estimator, which could contribute to good results. We consider their IAF results complementary to ours, where we work with moderate computational resources, and they consider the case of higher resource and power consumption. We also comment on batch size choices in Appendix B.3.
@@ 4.2.3 Varying properties of synthetic targets
i-ResNet $-0.21\pm 0.00$.19) $0.10\pm 0.00$.15) $-0.06\pm 0.00$.18) $0.20\pm 0.00$.27) $-0.00\pm 0.00$.10)
ResFlow $-0.37\pm 0.00$.21) $0.19\pm 0.00$.23) $-0.18\pm 0.00$.18) $0.07\pm 0.00$.25) $-0.09\pm 0.00$.11)
CNFEuler $0.30\pm 0.00$.28) $0.36\pm 0.00$.27) $-0.39\pm 0.00$.32) $-0.09\pm 0.00$.27) $0.01\pm 0.00$.15)
CNFRK $0.33\pm 0.00$.29) $-0.00\pm 0.00$.26) $-0.03\pm 0.00$.23) $-0.03\pm 0.00$.25) $0.07\pm 0.00$.13)
CNFRK(R) $0.02\pm 0.00$.29) $-0.12\pm 0.00$.38) $-0.43\pm 0.00$.27) $-0.38\pm 0.00$.32) $-0.24\pm 0.00$.15)
@@
IAF $0.50\pm 0.00$.25) $0.03\pm 0.00$.30) $-1.11\pm 0.00$.10) $-0.44\pm 0.00$.31) $-0.31\pm 0.00$.15)
IA-LR-NSF $-0.06\pm 0.00$.19) $0.10\pm 0.00$.14) $-0.65\pm 0.00$.13) $-0.88\pm 0.00$.24) $-0.41\pm 0.00$.11)
IA-RQ-NSF $-0.06\pm 0.00$.28) $0.00\pm 0.00$.24) $-0.93\pm 0.00$.11) $-0.67\pm 0.00$.32) $-0.51\pm 0.00$.14)
IA-NAFdeep $-0.16\pm 0.00$.20) $-0.05\pm 0.00$.44) $-0.83\pm 0.00$.18) $-0.84\pm 0.00$.20) $-0.51\pm 0.00$.13)
IA-NAFdense $-0.53\pm 0.00$.34) $-0.08\pm 0.00$.26) $0.34\pm 0.00$.22) $0.48\pm 0.00$.18) $0.09\pm 0.00$.14)
IA-NAFboth $-0.51\pm 0.00$.40) $-0.05\pm 0.00$.25) $0.80\pm 0.00$.20) $0.53\pm 0.00$.20) $0.27\pm 0.00$.15)
i-ResNet $-0.30\pm 0.00$.40) $-0.83\pm 0.00$.41) $-0.33\pm 0.00$.12) $-0.59\pm 0.00$.22) $-0.47\pm 0.00$.16)
ResFlow $-0.38\pm 0.00$.23) $-1.32\pm 0.00$.28) $-0.33\pm 0.00$.15) $-0.03\pm 0.00$.26) $-0.44\pm 0.00$.14)
Planar $0.43\pm 0.00$.19) $-0.64\pm 0.00$.19) $-0.02\pm 0.00$.14) $-0.19\pm 0.00$.22) $-0.07\pm 0.00$.10)
Radial $-0.11\pm 0.00$.51) $0.89\pm 0.00$.46) $1.65$ $1.65$ $1.12\pm 0.00$.18)
Sylvester $-0.40\pm 0.00$.15) $-0.31\pm 0.00$.44) $-0.16\pm 0.00$.08) $-0.32\pm 0.00$.23) $-0.29\pm 0.00$.11)

### v2

@@ 4.1.2 Best-case analysis for specific target families
Jump HMC ranks best on each family and is tied with Jump MH on non-Gaussian unimodal targets. In gradient-free sampling, Jump MH performs best on synthetic targets, while IMH is better on real-world targets. This is consistent with a sparse logistic regression analysis by [23], which found that adding global NF proposals to HMC outperforms classic HMC. They similarly found IMH to be better on a multimodal target compared to a sampler that cannot transition between modes. In our multimodal analyses, HMC ranks better than IMH on average, which could be due to the broad initialization of many chains and expressive HMC dynamics allowing chains to cross low target density barriers during the step size tuning phase.
[30] show that NeuTra HMC explores geometrically complex targets more efficiently than HMC. However, we find their ranks to be the same on non-Gaussian targets. Due to the high uncertainty in HMC, it is plausible that some NFs allow NeuTra HMC to rank better, though this is not the case in general. Moreover, NeuTra HMC is actually worse for Gaussian targets and does not improve results in general (see Figures 1(a) and 1(b)). This finding is complementary to the multivariate Gaussian target analysis by [23]. The authors found that NeuTra MCMC performs worse when the Gaussian NF is close to the target, but outperforms Jump MCMC if the NF fit is poor. This is consistent with the Gaussian targets in this benchmark being relatively simple to model, especially considering long NF training times and fairly large amounts of training data.
@@ 4.2.1 Jump performance on different target families
NICE $0.43\pm 0.00$.31) $1.45\pm 0.00$.14) $0.80\pm 0.00$.25) $0.00\pm 0.00$.32) $0.11\pm 0.00$.22)
Real NVP $0.80\pm 0.00$.25) $1.16\pm 0.00$.43) $0.29\pm 0.00$.36) $0.29\pm 0.00$.32) $0.11\pm 0.00$.22)
C-LR-NSF $0.14\pm 0.00$.72) $0.72\pm 0.00$.58) $0.51\pm 0.00$.30) $0.40\pm 0.00$.19) $0.24\pm 0.00$.20)
C-RQ-NSF $0.87\pm 0.00$.19) $1.01$ $0.51\pm 0.00$.38) $0.43\pm 0.00$.39) $0.00\pm 0.00$.24)
C-NAFdeep $0.58\pm 0.00$.19) $0.14\pm 0.00$.87) $0.72\pm 0.00$.35) $0.80\pm 0.00$.25) $0.63\pm 0.00$.17)
@@
i-ResNet $0.80\pm 0.00$.36) $0.14\pm 0.00$.58) $0.07\pm 0.00$.60) $0.98\pm 0.00$.25) $0.64\pm 0.00$.20)
ResFlow $0.22\pm 0.00$.62) $0.29\pm 0.00$.14) $0.80\pm 0.00$.18) $0.18\pm 0.00$.33) $0.11\pm 0.00$.21)
CNFEuler $0.14\pm 0.00$.35) $0.87\pm 0.00$.43) $0.36\pm 0.00$.62) $0.54\pm 0.00$.31) $0.39\pm 0.00$.21)
CNFRK $0.29\pm 0.00$.65) $0.00\pm 0.01$.59) $0.07\pm 0.00$.48) $0.72\pm 0.00$.27) $0.27\pm 0.00$.26)
CNFRK(R) $0.43\pm 0.00$.31) $1.16\pm 0.00$.14) $0.80\pm 0.00$.27) $0.54\pm 0.00$.32) $0.64\pm 0.00$.17)
@@
The best-performing architectures on the entire benchmark are CNFRK(R), i-ResNet, and CNFEuler. All obtain a good $\overline{r}$ on each family except CNFEuler on Gaussians. CNFRK(R) achieves better overall $\overline{r}$ than CNFRK, suggesting that regularization is beneficial for continuous NF models. NAF models rank the worst, which could be due to having a very high parameter count. A high parameter count increases the space of possible solutions, which can result in slower training. Additionally, NAF models use neural network transformers, which are more difficult to optimize than affine or spline maps. Residual NFs and continuous NFs have fewer restrictions on the form of their bijections than autoregressive NFs. This could explain their better performance on real-world targets, whose complex geometry requires expressive bijections to be modeled. Having observed how the bijection form can contribute to sampling quality, we also investigate the effect of the NF parameter count. In Figure 2, we plot the trainable parameter count against the SR of the NF.
@@
We find a noticeable improvement in performance with $10^{3}$ to $10^{4}$ parameters where we observe a dip in SR. The results verify our claim that one must choose hyperparameters that create a suitably expressive architecture. Furthermore, they suggest that NF parameter counts in this range are a suitable default choice. For SVI with Real NVP, [2] generally recommend using 10 or more layers for targets with complex geometry, as well as many hidden units for high-dimensional targets. Our analysis shows that these recommendations do not directly translate into NFMC, as the best results are generally achieved by NFs with few parameters.
@@ 4.2.2 Preconditioning quality on different target families
C-LR-NSF $0.62\pm 0.00$.42) $0.33\pm 0.00$.33) $0.37\pm 0.00$.49) $0.20\pm 0.00$.32) $0.10\pm 0.00$.21)
C-RQ-NSF $0.87\pm 0.00$.33) $0.33\pm 0.00$.17) $0.04\pm 0.00$.25) $0.45\pm 0.00$.33) $0.36\pm 0.00$.18)
C-NAFdeep $1.07\pm 0.00$.08) $0.50\pm 0.00$.17) $40\pm 0.00.11$) $0.62\pm 0.00$.41) $0.88\pm 0.00$.20)
C-NAFdense $0.33\pm 0.00$.45) $0.91\pm 0.00$.58) $1.11\pm 0.00$.23) $0.56\pm 0.00$.30) $0.47\pm 0.00$.22)
C-NAFboth $0.04\pm 0.00$.32) $0.99\pm 0.00$.33) $0.25\pm 0.00$.29) $0.08\pm 0.00$.37) $0.02\pm 0.00$.20)
IAF $0.54\pm 0.00$.44) $1.16\pm 0.00$.33) $0.04\pm 0.00$.60) $0.39\pm 0.00$.27) $0.41\pm 0.00$.21)
IA-LR-NSF $0.25\pm 0.00$.55) $0.25\pm 0.00$.08) $0.37\pm 0.00$.62) $0.07\pm 0.00$.41) $0.03\pm 0.00$.25)
IA-RQ-NSF $0.62\pm 0.00$.44) $0.50\pm 0.00$.33) $0.87\pm 0.00$.31) $0.30\pm 0.00$.28) $0.52\pm 0.00$.17)
IA-NAFdeep $0.83\pm 0.00$.24) $1.16\pm 0.00$.33) $0.37\pm 0.00$.40) $0.45\pm 0.00$.38) $0.59\pm 0.00$.20)
IA-NAFdense $0.54\pm 0.00$.27) $0.91\pm 0.00$.41) $0.87\pm 0.00$.37) $0.26\pm 0.00$.34) $0.53\pm 0.00$.19)
IA-NAFboth $0.29\pm 0.00$.25) $0.58\pm 0.00$.08) $0.12\pm 0.00$.46) $0.25\pm 0.00$.38) $0.07\pm 0.00$.20)
i-ResNet $0.58\pm 0.00$.27) $0.58\pm 0.00$.58) $0.70\pm 0.00$.18) $0.39\pm 0.00$.21) $0.52\pm 0.00$.13)
ResFlow $0.95\pm 0.00$.39) $1.07\pm 0.00$.08) $0.62\pm 0.00$.24) $0.43\pm 0.00$.22) $0.66\pm 0.00$.14)
Planar $0.29\pm 0.00$.55) $0.25\pm 0.00$.08) $0.04\pm 0.00$.21) $0.11\pm 0.00$.31) $0.04\pm 0.00$.18)
Radial $1.16\pm 0.00$.39) $1.32\pm 0.00$.33) $1.57\pm 0.00$.08) $1.07\pm 0.00$.21) $1.23\pm 0.00$.13)
Sylvester $0.74\pm 0.00$.20) $0.58\pm 0.01$.07) $0.00\pm 0.00$.24) $0.67\pm 0.00$.34) $0.47\pm 0.00$.19)
CNFEuler $1.49$ $0.99$ $0.00\pm 0.00$.58) $0.58\pm 0.00$.39) $0.70\pm 0.00$.24)
@@
As groups, both IA and residual NFs perform the best across all targets (last column). This is consistent with IA and radial flows having been designed to improve variational inference [48, 43, 33], which forms the crucial NF warm-up phase in NeuTra MCMC. Interestingly, the radial flow performs better than IA within NeuTra MCMC despite the fact that IAF has been proposed to address the limitations of the former. NeuTra MCMC can benefit from simpler preconditioning of the radial flow, which does not use neural networks, compared to more expressive bijections. The form of the bijection could also play a role, as the simple planar NF ranks worse than radial and Sylvester flows.
@@
Our results also shed light on [30], who precondition HMC with the IAF bijection, which improves moment estimates over HMC on the funnel target. However, the second column of Table 3 shows that IAF is one of the worst-performing NFs for the funnel and Rosenbrock distributions. The discrepancy could be due to differences in implementation, namely the number of chains: they use 16384 parallel chains on the GPU, whereas we use 100 on the CPU. We cannot afford such a big number of GPU chains, as it would require a prohibitive amount of computational resources for a fair comparison with other, usually slower NFs. [30] also use a bigger SVI batch size compared to our single-sample unbiased loss estimator, which could contribute to good results. We consider their IAF results complementary to ours, where we work with moderate computational resources, and they consider the case of higher resource and power consumption. We also comment on batch size choices in Appendix B.3.
@@ 4.2.3 Varying properties of synthetic targets
i-ResNet $0.21\pm 0.00$.19) $0.10\pm 0.00$.15) $0.06\pm 0.00$.18) $0.20\pm 0.00$.27) $0.00\pm 0.00$.10)
ResFlow $0.37\pm 0.00$.21) $0.19\pm 0.00$.23) $0.18\pm 0.00$.18) $0.07\pm 0.00$.25) $0.09\pm 0.00$.11)
CNFEuler $0.30\pm 0.00$.28) $0.36\pm 0.00$.27) $0.39\pm 0.00$.32) $0.09\pm 0.00$.27) $0.01\pm 0.00$.15)
CNFRK $0.33\pm 0.00$.29) $0.00\pm 0.00$.26) $0.03\pm 0.00$.23) $0.03\pm 0.00$.25) $0.07\pm 0.00$.13)
CNFRK(R) $0.02\pm 0.00$.29) $0.12\pm 0.00$.38) $0.43\pm 0.00$.27) $0.38\pm 0.00$.32) $0.24\pm 0.00$.15)
@@
IAF $0.50\pm 0.00$.25) $0.03\pm 0.00$.30) $1.11\pm 0.00$.10) $0.44\pm 0.00$.31) $0.31\pm 0.00$.15)
IA-LR-NSF $0.06\pm 0.00$.19) $0.10\pm 0.00$.14) $0.65\pm 0.00$.13) $0.88\pm 0.00$.24) $0.41\pm 0.00$.11)
IA-RQ-NSF $0.06\pm 0.00$.28) $0.00\pm 0.00$.24) $0.93\pm 0.00$.11) $0.67\pm 0.00$.32) $0.51\pm 0.00$.14)
IA-NAFdeep $0.16\pm 0.00$.20) $0.05\pm 0.00$.44) $0.83\pm 0.00$.18) $0.84\pm 0.00$.20) $0.51\pm 0.00$.13)
IA-NAFdense $0.53\pm 0.00$.34) $0.08\pm 0.00$.26) $0.34\pm 0.00$.22) $0.48\pm 0.00$.18) $0.09\pm 0.00$.14)
IA-NAFboth $0.51\pm 0.00$.40) $0.05\pm 0.00$.25) $0.80\pm 0.00$.20) $0.53\pm 0.00$.20) $0.27\pm 0.00$.15)
i-ResNet $0.30\pm 0.00$.40) $0.83\pm 0.00$.41) $0.33\pm 0.00$.12) $0.59\pm 0.00$.22) $0.47\pm 0.00$.16)
ResFlow $0.38\pm 0.00$.23) $1.32\pm 0.00$.28) $0.33\pm 0.00$.15) $0.03\pm 0.00$.26) $0.44\pm 0.00$.14)
Planar $0.43\pm 0.00$.19) $0.64\pm 0.00$.19) $0.02\pm 0.00$.14) $0.19\pm 0.00$.22) $0.07\pm 0.00$.10)
Radial $0.11\pm 0.00$.51) $0.89\pm 0.00$.46) $1.65$ $1.65$ $1.12\pm 0.00$.18)
Sylvester $0.40\pm 0.00$.15) $0.31\pm 0.00$.44) $0.16\pm 0.00$.08) $0.32\pm 0.00$.23) $0.29\pm 0.00$.11)